Suhami Chadhuri has a one-sixth interest in a partnership. She sold five-ninths of her interest for $2900. (a) What was the total amount of Ms. Chadhuris interest before selling? (b) What is the value of the partnership? (a) The amount is s (Round to the nearest cent as needed.) (b) The value is S (Round to the nearest cent as needed.)
Answer #1

(a) 5/9th of 1/6th of total partnership = $2900

Her interest before selling = 1/6th.

Thus 1/6th of total partnership = 2900*9/5

= $5,220.00

(b) Value of partnership = 5220*6

= $31,320.00